OpenStudy (anonymous):

solve x^4-12x-5=0 given that -1+2i is a root!

Parth (parthkohli):

-1 - 2i is also a root. Polynomial division now

OpenStudy (anonymous):

if \(-1+2i\) is a root then one of the factors is \(x^2+2x+5\)

OpenStudy (anonymous):

yeah! i know that -1-2i is also a root!

OpenStudy (anonymous):

how did you get that?! :o

OpenStudy (anonymous):

if that is not clear, we can work backwards \[x=-1+2i\] \[x+1=2i\] \[(x+1)^2=-4\] \[x^2+2x+5=0\]

Parth (parthkohli):

\[(-1 - 2i)(-1+2i)\times p(x) = x^4 - 12x - 5\]Your job is to find \(p(x)\), which'd turn out to be what sat said.

Parth (parthkohli):

But yeah, the guy is a genius. You should listen to him before doing polynomial division.

OpenStudy (anonymous):

or if you don't want to work backwards you can memorize that if \(a+bi\) is a root your quaddratic is \[x^2-2ax+a^2+b^2\]

Parth (parthkohli):

Omg, I messed up.

OpenStudy (anonymous):

okay i understand how you got x2+2x+5=0!

OpenStudy (anonymous):

That will not work @ParthKohli \(p(x) = \cfrac{x^4 - 12x - 5 }{5}\)

OpenStudy (anonymous):

or if you really want to be annoyed, multiply \[(x-(-1+2i))(x-(-1-2i))\] which is not as hard as it seems, although it is easier just to memorize it

Parth (parthkohli):

\[(x + 1 - 2i)(x + 1 + 2i)\]

OpenStudy (anonymous):

Yep, its now right.

OpenStudy (anonymous):

nah, leave in the parentheses

OpenStudy (anonymous):

yeah my teacher taught me (x−(−1+2i))(x−(−1−2i))

OpenStudy (anonymous):

yeah that works if you want to do it that way

OpenStudy (anonymous):

so um is x2+2x+5=0 the answer? or are theyre more steps?

OpenStudy (anonymous):

first outer inner last first is \(x^2\) last is \(1^2+2^2=5\) and outer, inner add up to \(-2x\)

Parth (parthkohli):

You should use that quadratic to get your solutions.

OpenStudy (anonymous):

how do i use this quadratic to get my solutions?

Parth (parthkohli):

Just get that quadratic's solutions!

OpenStudy (anonymous):

oops i meant "inner" adds up to \(2\) then divide or just think

OpenStudy (anonymous):

\((x-(-1 + 2i) )(x- (-1-2i) )\) \((x+1 - 2i )(x+1 + 2i) \) \(((x+1)-2i)((x+1) + 2i)\) Use (a+b)(a-b) identity now. Solution of quadratic equation : \(ax^2 + bx + c = 0 \) --> general form \(x = \cfrac{-b \pm \sqrt{b^2 - 4ac}}{2a} \)

OpenStudy (anonymous):

\[x^4-12x-5=(x^2+2x+5)(something)\]

OpenStudy (anonymous):

pretty clear that the first term of the "something" is \(x^2\) and the last term is \(-1\) just need the middle term

OpenStudy (anonymous):

sorry um, where did you get the x4−12x−5?

OpenStudy (anonymous):

i looked at the original question

OpenStudy (anonymous):

x2+2x+5=0 do i just plug this into the quadratic formual?

OpenStudy (anonymous):

no no

Parth (parthkohli):

Sorry, I messed up again :-P

Parth (parthkohli):

You should get that something and its solutions

OpenStudy (anonymous):

that is the one that has zeros of \(-1+2i\) and \(-1-2i\) we already know those zeros

OpenStudy (anonymous):

right! so i divide x4−12x−5 by (x2+2x+5)

OpenStudy (anonymous):

you want the other two zeros, so you have to factor \[x^4-12x-5\]

OpenStudy (anonymous):

you can do that if you want a pain

OpenStudy (anonymous):

i dont want a pain :(

OpenStudy (anonymous):

i would do what i said above, write \[x^4-12x-5=(x^2+2x+5)(something)\] and think about what the something is

OpenStudy (anonymous):

it is easier than dividing by long division you know it has to look like \[x^4-12x-5=(x^2+2x+5)(x^2+bx-1)\] the only think you don't know is \(b\)

OpenStudy (anonymous):

i hope that is clear the constant is \(-5\) and you have a \(5\) so the constant must be \(-1\) for the second factor

OpenStudy (anonymous):

okay in the text book it says that SOMETHING is x^2-2x-1

OpenStudy (anonymous):

fine, we can find the \(-2\) now

OpenStudy (anonymous):

the \(x^2\) part is obvious i hope, as is the \(-1\) part

OpenStudy (anonymous):

yes i understand the -1 and the x!

OpenStudy (anonymous):

for example, when you multiply out in what i wrote you will get \(-2x+5bx\) and you know the answer has to be \(-12x\) so if \(-2x+5bx=-12x\) then \(-2+5b=-12\) and so \(b=-2\) that is one way

OpenStudy (anonymous):

or another way: if you multiply you will get \(bx^2+2x^3\) but you know you have no \(x^3\) and so \(b=-2\)

OpenStudy (anonymous):

typo there, i meant \(bx^3+2x^3=0\) and so \(b=-2\)

OpenStudy (anonymous):

in any case you must have \(b=-2\) making it \[x^4-12x-5=(x^2+2x+5)(x^2-2x-1)\]

OpenStudy (anonymous):

or if thinking is too annoying, you can always use long division which i think it totally annoying

OpenStudy (anonymous):

i think i understand your way a bit, but just not very well. I probably will try your way and use long division to check

OpenStudy (anonymous):

Thanks a lot Satellite!